# Visualizationary: Automating Design Feedback for Visualization Designers using LLMs

### v1

@@ Abstract
Interactive visualization editors empower people to author visualizations without writing code, but do not guide them in the art and craft of effective visual communication. In this paper, we explore the potential for using an off-the-shelf Large Language Model (LLM) to provide actionable and customized feedback to visualization designers. Our implementation, called Visualizationary, showcases how ChatGPT can be used in this manner using two components: a preamble of visualization design guidelines and a suite of perceptual filters extracting salient metrics from a visualization image. We present findings from a longitudinal user study involving 13 visualization designers—6 novices, 4 intermediate ones, and 3 experts—authoring a new visualization from scratch over the course of several days. Our results indicate that providing guidance in natural language using an LLM can aid even seasoned designers in refining their visualizations. All supplemental materials accompanying this paper are available at https://osf.io/v7hu8.
@@
Visualization, design, design critique, feedback, human-centered AI, large language models.
@@ 1 Introduction
Figure 1: System overview. Example of an interactive report on a visualization artifact generated using our novel analyze-clarify-guide-track (ACGT) framework. The report is generated by automatically analyzing (“analyze”) the artifact using an expandable collection of image filters grouped into categories: salience, text, visual representation, color, and accessibility. After reporting results from each image filter (center column), the report then interprets (“clarify”) the findings using natural language understandable to a novice visualization designer. This is followed by natural language suggestions (“guide”) on how to address weaknesses in the design. Finally, the right side of the report shows changes over time (“track”) as the designer iteratively refines the artifact.
To democratize the transformative power of data in today’s information society, it is not sufficient that people get access to visualizations created by others; they must also be empowered to author their own visualizations. As a result, recent years have seen an influx of interactive tools for creating visualizations without writing code, such as iVisDesigner [54], Data Illustrator [39] and Data Animator [64], Charticulator [55], and Lyra [59] and Lyra 2 [74]. However, no matter how advanced these tools become, a fundamental barrier remains: authoring effective visualizations is a complex task requiring visual design skills, aesthetics, and experience that not all would-be designers possess.
Fortunately, visualization designers have nothing to lose but their chains. We propose Visualizationary, a tool to seize the means of visualization design by using off-the-shelf Large Language Models (LLMs) such as GPT or LLaMA. The basic idea behind Visualizationary is to investigate how an LLM can be used to aid the iterative design process of a visualization artifact in a workflow that we call analyze-clarify-guide-track (ACGT):
@@
We report on a web-based implementation of the Visualizationary system using the OpenAI GPT API, showcasing how a “vanilla” LLM can be utilized for this highly complex visualization design task. In particular, we translate chart images into text using a VLM (vision-language model) [37] and then generate automatic design guidance using an LLM as a template and text processor. The tool presents the guidance in an interactive report organized into a hierarchical structure that enables the designer to drill into the findings to the desired level of detail. We also present findings from a longitudinal study conducted with 6 novice, 4 intermediate, and 3 expert visualization designers over a period spanning a few days. During the study, each designer worked with Visualizationary to create a new visualization from scratch. We report on four successive revisions of their visualization as well as their quantitative and qualitative feedback. While Visualizationary and the ACGT workflow is primarily designed for novices, our findings show how the approach can be useful even for an experienced designer by summarizing feedback in an easily navigated interactive report format.
@@ 2 Background
Here we present the state of the art in three topics: design feedback, automating visualization design feedback, and LLMs.
@@ 2.1 Design Feedback
Feedback is essential in all design disciplines, and it typically comes from two sources: user and usage feedback versus peer and supervisor feedback. In academic contexts, the former, which emphasizes validation through empirical evaluation, is more prevalent. Munzner’s nested model for visual design and validation [49] primarily incorporates this type of feedback. Conversely, peer feedback holds greater prevalence in industry and design practice, particularly for designs aimed at widespread use. Nevertheless, this methodology has also permeated academic circles, notably within interaction design [5, 4].
For feedback to be effective, it must not only be understandable by the intended recipient, but it must also provide enough information so that the recipient can act upon the feedback (if necessary). What Munzner calls low-level feedback [50] is focused on fine-grained insights and improvements, and is often actionable because of its low abstraction level: e.g., “this red color is a poor choice for a color-blind person,” “this bar must be made larger,” and “this text uses a font with poor legibility.” However, higher-level design feedback in the form of critique, qualitative evaluation, and abstract metrics can be much harder to understand, let alone act upon. For example, learning that your visualization suffers from “poor contrast,” “poor color choices,” or is “not aesthetically pleasing” can be challenging to address.
@@ 2.3 Large Language Models
A language model is a statistical model that predicts the likelihood of word sequences in natural language based on a large corpus of textual training data. Recent work [8, 69, 24] shows that the primary discriminator in accomplishing complex tasks using language models is their scale. In light of this trend, there has been a growing number of works on so-called “large” language models (LLMs) [8, 51, 14, 65, 2]. LLMs have trillions of model parameters, and thus, training these models is computationally demanding. It needs a large training dataset, GPU hardware acceleration, and a machine equipped with them. This means that training an LLM is expensive in time and resources.
A standard paradigm for addressing the training costs is to pre-train LLMs on a large corpus of datasets and then allow users to fine-tune these models on a set of desired tasks, e.g., question answering, language translation, and text summarization [8, 17, 53, 2, 36, 26]. In pre-training, a model is typically trained to do next-word prediction, i.e., given a prompt, the model returns a sequence of word tokens likely to complete the prompt. This process is performed by third-parties such as OpenAI, Google, or Meta, and such pre-trained models are offered to users as a form of internet service, e.g., ChatGPT [51], or as-is, e.g., open-source models such as Vicuna [14] or LLaMA [65]. In particular, services such as ChatGPT offer a personalized experience by separating sessions for each user and fine-tuning the pre-trained model based on the user’s inputs. Recent models further support multi-modality, with images and text being combined, and enable them to answer questions, e.g., in data visualizations: “How much does X outperform Y?” [37, 68]
Our work leverages both types of pre-trained LLMs to generate feedback: (1) Multi-modal LLMs that take a user’s data visualization and questions on the visualization as input and return the answers to the questions, and (2) uni-modal LLMs (or standard ones) that receive text inputs and return the corresponding outputs. We use multi-model LLMs to generate textual descriptions for a user’s visualization, and we then combine the descriptions and the queries as the input to uni-modal LLMs, asking the model answer queries as design feedback.
@@ 3.1 Target Audience
Our work in this paper primarily targets novice or intermediate designers who do not have the required experience or expertise to critique their own work with an objective eye, or the ability to make the required changes to address problems. In other words, our automated design feedback mechanism is not intended to be a replacement for a trained designer’s critical eye. However, because design critique is critical during any design process, automated design feedback can be useful even for experts [61]. Similarly, automation is no replacement for a human peer or mentor to give feedback and suggestions to the novice designer. However, mentors or peers may not always be available.
3.2 Rationale
Based on Munzner’s visualization design and nested evaluation frameworks [49, 50], interaction design critique [4], and guidelines for visualization design [23, 47, 55], we identify several benefits for automating visualization design feedback:
  • Efficiency: An automated mechanism can provide immediate feedback to support rapid design iterations.
  • Consistency: Automated feedback systems apply consistent evaluation criteria, reducing subjectivity and bias as well as human error in assessments. Human feedback, especially if collected from different sources, may be less consistent and sometimes even conflicting.
  • Scalability: Compared to a human evaluator, an automated design feedback mechanism can handle a large volume of evaluations simultaneously, making it feasible to assess multiple design variations.
@@ 3.2 Rationale
  • Real-time: An automated visualization feedback system can provide data in real-time, enabling designers to make immediate improvements during the design process.
@@ 3.3 Analyze-Clarify-Guide-Track (ACGT) Workflow
    Analyze: To provide written feedback to the designer, we first need to collect textual data from the visualization screenshot; i.e., we must make the transition from images to text. We do this using automated vision filters, such as color contrast, visual hierarchy, and layout consistency (see Sec. 3.4). The outcome of this analysis is a report with metrics drawn from the visualization artifact.
@@
    Clarify: Interpreting visualization metrics on their own can be challenging [61], let alone to use as a basis for revisions. The clarify phase translates results from the perceptual filters into understandable terms, thus bridging the gap between technical analysis and practical design improvements. It breaks down complex issues identified in the analysis into layman’s terms, explaining why certain changes are needed and what needs attention.
@@
    Guide: It is not enough to know the problem; we must also know how to fix it. Our mechanism provides such practical guidance to the designer as natural language on how to address the concerns identified during the analysis. This guidance may include recommendations, actionable insights, or best practices for improvements.
@@
3.4 Visualization Filters
We use visualization filters to extract perceptual characteristics from visualizations, with the goal of providing pragmatic and actionable design feedback. These filters are a suite of computer vision and vision-language models (VLMs) that we collectively call visualization filters. Note that the list is not exhaustive.
[…]
    Virtual Eyetracker refers to the extent to which areas within a scene capture the attention of an observer [58].
[…]
    Textual characters are visual objects that are used for labels, titles, and legends in visualization [62].
[…]
    Visual representations can generate feedback specific to the technique, such as visualization linting [13, 30].
[…]
    Color perception is a basic building block of data visualization [50, 57]. Visualization practice stipulates using a limited number of distinguishable and easily named colors [29], potentially as a function of the mark used [63].
[…]
    Accessibility is emerging as an important new topic in the visualization community [15, 18, 44, 20]. Automated feedback can be highly useful even for seasoned designers.
LLMs are capable of generating human-like natural language responses under various tasks. To that end, we first extract numerical metrics from a visualization and pass it to an LLM. We also instill design knowledge in the LLM so that it can guide designers. We explain this idea in the next section.
4 The Visualizationary Tool
Visualizationary is a web-based visualization design feedback system that automates natural language critique for visualization designers. In this section, we describe its system architecture, overall workflow, and core components. We also describe the prompts we define for generating quality feedback from an LLM.
@@ 4.1 System Overview
Visualizationary is a web-based client-server system that tracks a visualization artifact throughout its design process. The workflow (Fig. 2) is based on the user—a visualization designer—iteratively uploading a screenshot of the current state of their visualization artifact and receiving automated feedback from the system. The feedback then presumably helps the designer make improvements to the artifact in the next design iteration. However, unlike existing visualization design feedback tools such as Perceptual Pat [61], which merely report metrics, Visualizationary implements the ACGT workflow from Sec. 3: beyond analyzing the image, it also explains (“clarify”) the feedback using natural language, suggests ways for addressing shortcomings (“guide”), and tracks the changes to the design over multiple iterations.
@@
This workflow is realized through five components: (1) a web-based visualization design management interface, (2) interactive design reports providing feedback at user-controlled levels of detail, (3) a revision tracker, (4) a suite of computer vision filters for extracting visualization data into symbolic form, and (5) a design guidance tool that uses an off-the-shelf Large Language Model (LLM) with a prompt preamble to explain the feedback and guide the user. The former two belong to Visualizationary’s frontend, whereas the remaining three are part of the backend. Below we describe each of these components.
4.2 Visualization Design Management Interface
@@
The frontend for Visualizationary is a web-based single-page application for managing the iterative design process of a visualization artifact (Fig. 3). A project is a container for a design process, including all of the visualization revisions and their corresponding design reports. The frontend has operations for creating, editing, and deleting projects associated with a user.
Figure 3: Visualizationary design management interface. The system shows an overview of the timeline of revisions for a visualization artifact. Each revision consists of a screenshot of the artifact as well as an interactive design report. The right column shows the revision tracker, which provides a high-level view of the overall design trajectory.
@@
Viewing a project shows a timeline of the revisions and allows the user to bring up the details for each. Projects can be extended by uploading a new screenshot of the visualization artifact in .jpg or .png file (see Fig. 1) and then launching the analyze-clarify-guide-track workflow on the new revision. This process typically takes a couple of minutes to complete, after which a new design report is added to the timeline for the designer’s perusal. In this way, the Visualizationary interface is designed to manage the entire design process of a visualization artifact, from early sketches to final production.
4.3 Visualization Metrics
We present metrics to quantitatively assess the visualization image using visualization filters. To each evaluation category, we present the goal, metric used, and decisions. The decision is made based on predefined heuristics about the metric and also on the judgment of LLMs, backed by the design knowledge added as a preamble. This heuristic is described in our supplementary material. Again, note that the list is not exhaustive and can be extended.

### v3

@@ Abstract
Interactive visualization editors empower users to author visualizations without writing code, but do not provide guidance on the art and craft of effective visual communication. In this paper, we explore the potential of using an off-the-shelf large language models (LLMs) to provide actionable and customized feedback to visualization designers. Our implementation, Visualizationary, demonstrates how ChatGPT can be used for this purpose through two key components: a preamble of visualization design guidelines and a suite of perceptual filters that extract salient metrics from a visualization image. We present findings from a longitudinal user study involving 13 visualization designers—6 novices, 4 intermediates, and 3 experts—who authored a new visualization from scratch over several days. Our results indicate that providing guidance in natural language via an LLM can aid even seasoned designers in refining their visualizations. All our supplemental materials are available at https://osf.io/v7hu8.
@@
Visualization design, Design critique, Feedback, Human-centered AI, large language models.
@@ 1 Introduction
To democratize the transformative power of data in today’s information society, it is not sufficient that people get access to visualizations created by others; they must also be empowered to author their own visualizations. In response, recent years have seen an influx of interactive tools that enable users to create visualizations without writing code, such as iVisDesigner [51], Data Illustrator [38], and Lyra [56]. However, no matter how advanced these tools become, a fundamental barrier remains: authoring effective visualizations is a complex task requiring visual design skills, an understanding of aesthetics, and experience—expertise that not all would-be designers possess.
Fortunately, visualization designers have nothing to lose but their chains. We introduce Visualizationary, a system to seize the means of visualization design by using off-the-shelf large-language models (LLMs). Our system enhances the design of communicative visualizations by providing perceptual feedback. While it may not fully replace feedback from peers or colleagues, it serves as a valuable alternative when it is impossible to gain human feedback and can benefit designers of all expertise levels, from novices to seasoned ones.
Although competing approaches for improving visualization designs, such as heuristics (like small, smart defaults in Lyra) and Grammar of Graphics (GoG)-style specifications have their own advantages, they also carry limitations due to their deterministic nature. Heuristic-driven methods lack human intervention, which can constrain creativity and limit the potential for novel ideas. GoG-style specifications, despite offering precise chart descriptions, lead to producing standardized solutions, reducing the diversity of possible visualization outcomes.
In contrast, Visualizationary automates perceptual feedback using a suite of automated techniques that simulate how a viewer would perceive a visualization. By providing actionable insights without the need to standardize changes, Visualizationary preserves the designer’s creativity while enhancing effectiveness.
The basic idea behind Visualizationary is to investigate how an LLM can be used to aid the iterative design process of a visualization artifact in a workflow that we call analyze-clarify-guide-track (ACGT):
@@
We present a web-based implementation of Visualizationary that leverages a server-side LLM, demonstrating how a “vanilla” LLM can be used for this highly complex visualization design task. Visualizationary operates as follows: We first translate chart images into text using a vision-language model (VLM) [36] and then generate automatic design guidance using an LLM as a template and text processor. Users can also upload their existing designs as an image file, leaving them free to use a tool of their own choice. The system presents the guidance in an interactive report organized hierarchically, enabling designer to drill into the findings at their desired level of detail.
In evaluation, we present findings from a longitudinal study conducted with 6 novice, 4 intermediate, and 3 expert visualization designers over a period spanning a few days. During the study, each designer worked with Visualizationary to create visualizations and improve their designs. We report on four successive revisions of their visualization as well as their quantitative and qualitative feedback from 13 visualization designers of varying expertise, from a novice with only 1 year of experience in the field to an expert with 11 years of experience. The designs created by these designers are then assessed by 3 senior visualization researchers. Our findings show the effectiveness as well as shortcomings of our concept from various perspectives, including level of expertise, type of feedback, and design optimization. These insights contribute to a deeper understanding of what constitutes effective design feedback for visualizations.
@@ 2 Background
Here we provide a brief overview of three key areas, required for understanding our work: design feedback, automating visualization design feedback, and large-language models (LLMs).
@@ 2.1 Design Feedback
Feedback is essential in all design disciplines, and it typically comes from two sources: user and usage feedback versus peer and supervisor feedback. In academic contexts, the former, which emphasizes validation through empirical evaluation, is more prevalent [47, 74, 75, 50, 34]. Munzner’s nested model for visual design and validation [47] primarily incorporates this type of feedback. Conversely, peer feedback holds greater prevalence in practice, particularly for designs aimed at widespread use [1, 49, 59]. Nevertheless, this methodology has also permeated academic circles, notably within interaction design [5, 4, 14].
For feedback to be effective, it must not only be understandable by the intended recipient, but it must also provide enough information so that the recipient can act upon the feedback (if necessary). What Munzner calls low-level feedback [48] is focused on fine-grained insights and improvements, and is often actionable because of its low abstraction level: e.g., “this red color is a poor choice for a color-blind person,” “this bar must be made larger,” and “this text uses a font with poor legibility.” However, higher-level design feedback in the form of critique, qualitative evaluation [12], and abstract metrics can be much harder to understand, let alone act upon. For example, learning that your visualization suffers from “poor contrast,” “poor color choices,” or is “not aesthetically pleasing” can be challenging to address.
@@ 2.3 Large Language Models
A language model is a statistical model that predicts the likelihood of word sequences in natural language based on a large corpus of textual training data. Recent work [8, 70, 26] shows that the primary discriminator in accomplishing complex tasks using language models is their scale. In light of this trend, there has been a growing number of works on so-called “large” language models (LLMs). LLMs have billions of model parameters, and thus, training these models is computationally demanding.
A standard paradigm for addressing the training costs is to pre-train LLMs on a large text corpus and then allow users to fine-tune these models on a set of desired tasks, e.g., question answering, language translation, and text summarization. In pre-training, a model is typically trained to do next-word prediction, i.e., given a prompt, the model returns a sequence of word tokens likely to complete the prompt. This process is performed by third-parties such as OpenAI, Google, or Meta. In particular, services such as ChatGPT offer a personalized experience by separating sessions for each user and fine-tuning the pre-trained model based on the user’s inputs. Recent models further support multi-modality, with images and text being combined, and enable them to answer questions, e.g., in data visualizations: “How much does X outperform Y?” [36, 67]
Our work leverages both types of pre-trained LLMs to generate feedback: (1) Multi-modal LLMs that take a user’s data visualization and questions on the visualization as input and return the answers to the questions, and (2) uni-modal LLMs (or standard ones) that receive text inputs and return the corresponding outputs. We use multi-model LLMs to generate textual descriptions of a user’s visualization. These descriptions, along with user queries, are then fed into uni-modal LLMs, which generate design feedback based on the provided input. To minimize hallucination, we provide strict guidance as a preamble so that the language model does not provide random and inconsistent feedback [44].
@@ 3.1 Target Audience
Our work in this paper primarily targets novice or intermediate designers who do not have the required experience or expertise to critique their own work with an objective eye, or the ability to make the required changes to address problems. In other words, our automated design feedback mechanism is not intended to be a replacement for a trained designer’s critical eye. Because design critique is critical during any design process, we even find that automated design feedback is useful even for experts (see §6.2).
3.2 Design Rationale
Based on Munzner’s visualization design and nested evaluation frameworks [47, 48], interaction design critique [4], and guidelines for visualization design [25, 45, 52], we identify several benefits for automating visualization design feedback. We explain our rationale by comparing our system’s conceptual approach with the type of feedback a human designer would provide:
  • R1
    Ubiquitous: An automated system can provide immediate feedback, enabling rapid design iterations anytime [22, 20] the designer needs support. By contrast, human feedback may not always be readily-available. Furthermore, the system should not depend on the authoring tool in use.
  • R2
    Real-time: Hence, an automated visualization feedback system can provide data in a real-time, enabling designers to make immediate improvements during the design process.
  • R3
    Consistency: Automated feedback systems apply consistent evaluation criteria, reducing subjectivity and bias as well as human error in assessments. Human feedback may be less consistent and sometimes even conflicting.
  • R4
    Scalability: Compared to a human evaluator, an automated design feedback mechanism can handle a large volume of evaluations for multiple users simultaneously.
  • R5
@@ 3.3 Analyze-Clarify-Guide-Track (ACGT) Workflow
    Analyze: To provide written feedback to the designer, we first need to collect textual data from the visualization screenshot; i.e., we must make the transition from images to text. We do this using automated vision filters, such as color contrast, visual hierarchy, and layout consistency (see § 4.4). The outcome of this analysis is a report on the visualization artifact.
@@
    Clarify: Interpreting visualization metrics on their own can be challenging [61], let alone to use as a basis for revisions. The clarify phase translates results from the perceptual filters into understandable terms (R3), thus bridging the gap between technical analysis and practical design improvements. It breaks down complex issues identified in the analysis into layman’s terms, explaining which and why certain changes are needed.
@@
    Guide: It is not enough to know the problem; we must also know how to fix it. Our mechanism provides such practical guidance to the designer as natural language on how to address the identified concerns. This guidance may include recommendations, actionable insights, or best practices.
@@
4 How Visualizationary Works
Visualizationary is an automated design feedback system that provides design critiques for visualization designers in natural language. The system is implemented as a web-based platform, where the front-end offers an interactive interface for users to explore various visual feedback, while the back-end supports these processes using off-the-shelf language models. Here we describe its architecture and core components in detail.
@@ 4.1 System Overview
Figure 2: Visualizationary interface. Example of an interactive report on a visualization artifact generated using our novel analyze-clarify-guide-track (ACGT) framework. The image is a depiction of our system, Visualizationary where a new chart image (Figure 1(B)) is being updated while the current report is about the past visualization image. The report is generated by automatically analyzing (“analyze”) the artifact using an expandable collection of image filters grouped into categories: salience, text, visual representation, color, and accessibility. After reporting results from each image filter (center column), the report then interprets (“clarify”) the findings using natural language understandable to a novice visualization designer. This is followed by natural language suggestions (“guide”) on how to address weaknesses in the design. Finally, the right side of the report shows changes over time (“track”) as the designer iteratively refines the artifact.
@@
Visualizationary is a web-based client-server system (R4) that tracks a visualization artifact throughout its design process. Our system can support multiple users at once, so we placed no time constraints on image uploads (R4). However, no instances occurred where multiple users accessed it simultaneously. The workflow (Fig. 1) is based on the user—a visualization designer—iteratively uploading a screenshot of the current state of their visualization artifact and receiving automated feedback from the system. The feedback then presumably helps the designer make improvements to the artifact in the next design iteration. However, unlike existing visualization design feedback systems such as Perceptual Pat [61], which merely report metrics, Visualizationary implements the ACGT workflow from §3: beyond analyzing the image, it also explains (“clarify”) the feedback using natural language (R3), suggests ways for addressing shortcomings (“guide”), and tracks the changes to the design over multiple iterations.
@@
This workflow is realized through three components: (1) a web-based visualization design interface (R1, R4, R5) (§ 4.2), (2) interactive design reports providing feedback at user-controlled levels of detail (§ 4.3), and (3) an automated feedback system for extracting visualization data and transforming into textual feedback (R5) (§ 4.4). The former two belong to Visualizationary’s frontend, whereas the third is a part of the backend. Afterwards, we detail our prompting methodologies for providing visualization feedback (§ 4.5), and our implementation notes (§ 4.6).
4.2 Visualizationary Design Management Interface
@@
The frontend for Visualizationary is a web-based single-page application for managing the iterative design process of a visualization artifact. A project is a container for a design process, including all of the visualization revisions and their corresponding design reports (see § 4.3). The frontend has operations for creating, editing, and deleting projects associated with a user.
@@
Viewing a project shows a timeline of the revisions and allows the user to bring up the details for each. Projects can be extended by uploading a new screenshot of the visualization artifact in .jpg or .png file (see Fig. 2) (R1) and then launching the analyze-clarify-guide-track workflow on the new revision. This process typically takes a couple of minutes to complete (R2), after which a new design report is added to the timeline for the designer’s perusal. In this way, the Visualizationary interface is designed to help the design process of a visualization artifact.
4.3 Visualization Design Reports
Results from the feedback backend are presented as interactive visualization design reports. A design report is an interactive document consisting of a hierarchy of expandable sections designed in a bottom-up fashion. Sections are organized based on the main categories of visualization filters in § 4.4. The contents of a section in the report include both natural language and illustrative images (such as heatmaps, extracted shapes, or gaze maps) (see Fig. 2 (C)). Each section provides a short overall summary of suggestions (Fig. 2 (C) ②) of the underlying data for that metric followed by two standardized paragraphs:
  • Explanations: Explanation of visualization feedback in natural language (“clarify” in the ACGT workflow) (e.g., Fig. 2 (C) ④).
  • Suggestions: Recommendations addressing shortcomings in the design for that specific metric (“guide”) (e.g., Fig. 2 (C) ③).
Visualizationary provides additional functionality. First, to help navigating the report, we indicate sections that require modification using a circular visual mark, either yellow or red next to the title of each section (see Fig. 2 (C) ①). If there exist at most one issue in a section, then the corresponding circle is yellow. If there is more than one, then the circle is orange. Second, to help designers quickly skim section summaries, we provide a short overview summary of responses per category in Fig. 2 (E).
We also aim to maintain transparency in the feedback process, given that both filters and LLMs can yield misleading feedback. By being transparent, designers can better determine whether to trust the model. For instance, the report shows how the system extracts chart data and calculates its metrics.
4.4 Feedback System
Here we present our feedback system (see Fig. 1). The feedback is provided using the ACGT Workflow, as explained in § 3.3. We use automated models to extract perceptual characteristics from visualizations, with the goal of providing pragmatic and actionable design feedback. These functions are a suite of computer vision and vision-language models (VLMs) that we collectively call visualization filters. We present a total of 10 filters from 5 topics. We first detail how these visualization filters are translated to design feedback (what corresponds to ACG of our ACGT workflow). Afterwards, we describe how the revision Tracker works.
Visualization Filters. We explain the filters according to the ACGT framework. To each filter, we provide the intended role, Analysis process, Clarification criteria, and Guidance to users. In a nutshell, the filters automatically analyze the visualizations. Then, using the information from these filters, clarification is made based on predefined heuristics about the metric and also on the judgment of LLMs, backed by the design knowledge added as a preamble. Finally, based on the clarification, the LLM provides actionable guidance to users in the form of textual feedback. We instill the clarifications along with the guidance relevant to the decisions into ChatGPT as a prompt (see § 4.5). Then, based on a series of template questions, ChatGPT provides the appropriate answer.
However, despite our intents, not all filters we provide are translated into actionable feedback. This heuristic as well as examples of the feedback are described in our supplementary material. Note that the list is not exhaustive and can be extended.